# Investigating a Model-Agnostic and Imputation-Free Approach for Irregularly-Sampled Multivariate Time-Series Modeling

### v2

@@ 3.2 Learning Embeddings for Time-Series with Missing Features
Figure 4: Multiple Time-series Baselines. Performance comparison between MCAR and Periodic masking with multiple TS Baselines imputed with SAITS. TS Baselines considered: Autoformer (Wu et al. 2021), PatchTST (Nie et al. 2022b), iTransformer (Liu et al. 2023), DLinear (Zeng et al. 2023), SimMTM (Dong et al. 2024).
3.3 Missing Feature-Aware Attention (MFAA)
The MFAA Layer illustrated in Figure 3 leverage the power of “masked-attention” for learning latent representations at every time-step using partially observed features. MFAA works by computing attention scores based on the partially observed features at a time-step $t$, which are then used to perform a weighted sum of observed features to obtain the latent representation $\mathbf{L}_{t}$. As shown in Figure 3, these latent representations are projected back using a linear layer, to the original input shape before being fed into the downstream model (here, the encoder-decoder based self-supervised learning framework). MFAA performs a masked cross-attention using a learnable query vector and observed data as keys and values. This separation of roles is inspired by similar architectures in multi-modal grounding, for example, in (Carion et al. 2020), where learnable object queries serve as abstract object representations to focus on distinct objects in an image without requiring predefined region proposals, enabling set-based prediction. Similarly, in our setting, the learnable queries capture the interactions among variates independent of time, enabling the model to attend to the most informative aspects of observed variates at any time-step fed through keys and values. This intuition aligns with the query-based mechanism in mTAN (Shukla and Marlin 2021), which introduces a structured way to aggregate information over observed time-series data. However, while mTAN uses discrete reference points on a fixed temporal grid to achieve this, our single learnable query generalizes across variates at every time step, allowing for a more flexible representation of feature interactions
Mathematical Formulation: To obtain attention scores from partially-observed features at a time-step, we apply a masked scaled-dot product operation followed by a softmax operation described as follows. We first define a learnable query vector $\mathbf{Q}\in\mathbb{R}^{1\times D}$ which is independent of the variates and time-steps. The positionally-encoded embeddings at time-step $t$, $\mathbf{Z}_{(t,:)}$, are used as key and value inputs in the MFAA Layer. Specifically, The query, key, and value vectors are defined using linear projections as follows:
$\hat{\mathbf{Q}}=\mathbf{Q}\mathbf{W^{Q}},\quad\hat{\mathbf{K}}_{t}=\mathbf{Z}_{(t,:)}\mathbf{W^{K}},\quad\hat{\mathbf{V}}_{t}=\mathbf{Z}_{(t,:)}\mathbf{W^{V}},$ (4)
Here, $\hat{\mathbf{Q}}\in\mathbb{R}^{1\times d_{k}}$ and $\hat{\mathbf{K}}_{t},\hat{\mathbf{V}}_{t}\in\mathbb{R}^{N\times d_{k}}$, where $d_{k}$ is the dimension of the vectors after linear projection. The linear projection matrices for the query, key, and values are defined as: $\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{D\times d_{k}}$ respectively. Note that the key $\hat{\mathbf{K}}_{t}$ and value $\hat{\mathbf{V}}_{t}$ vectors depend on the time-step $t$, while the query vector doesn’t change with time. We then define the Missing Feature-Aware Attention (MFAA) Score at a given time-step $t$ as a masked scalar dot-product of the query and key vector followed by normalization of the scores using a Softmax operation, formally defined as follows:
where $\mathbf{A}_{t}\in\mathbb{R}^{N}$ is the MFAA Score vector of size $N$ corresponding to the $N$ variates, and $\eta\rightarrow$-$\infty$ is a large negative bias. The negative bias term $\eta$ forces the masked-elements that correspond to the missing variates in the time-series to have an attention score of zero. Thus, by definition, the $i$-th element of the MFAA Score $\mathbf{A}_{(t,i)}\neq 0\implies\mathcal{M}_{(t,:)}\neq 0$. We compute the latent representation $\mathbf{L}_{t}$ as a weighted sum of the MFAA score $\mathbf{A}_{t}$ and the Value vector $\hat{\mathbf{V}}_{t}$ as follows:
$=\mathrm{Concat}(\mathbf{L}^{0}_{t},\mathbf{L}^{1}_{t},\dots,\mathbf{L}^{h}_{t})\cdot\mathbf{W}^{O}$
3.4 Putting Everything Together: Plugging MissTSM with any MTS Model
Figure 3 shows the overall framework of a Masked Auto-Encoder (MAE) (He et al. 2022) based time-series model integrated with MissTSM. For an input time-series $\mathbf{X}$, we apply the TFI embedding layer followed by the MFAA layer to learn a latent representation for every time-step. The latent representations are then used as input to the downstream model. In this work, we opted for a MAE-based time-series model as the default downstream or base model, primarily due to its recent success in time-series modeling and its ability to perform both time-series forecasting and classification tasks. Furthermore, out of the several state-of-the-art masked time-series modeling techniques, we intentionally chose the simplest variation of MAE, namely Ti-MAE (Li et al. 2023), to highlight the effectiveness of TFI and MFAA layers in handling missing values.
Figure 5: Classification F1 scores on three datasetsL EMG, Epilepsy, and Gesture. Masking fractions considered: 0.2, 0.4, 0.6, 0.8.
4 Experimental Setup
Baselines: We considered the following MTS models for forecasting: SimMTM (Dong et al. 2024), PatchTST (Nie et al. 2022b), AutoFormer (Wu et al. 2021), DLinear (Zeng et al. 2023), and iTransformer (Liu et al. 2023). Four imputation techniques were considered: a $2^{\text{nd}}$-order spline imputation (McKinley and Levine 1998), a simple k-Nearest Neighbor approach (Tan et al. 2019), SAITS (Du, Côté, and Liu 2023), and BRITS (Cao et al. 2018). We considered the following imputation-free IMTS models: GRU-D (Che et al. 2018), Latent ODE (Rubanova, Chen, and Duvenaud 2019). Please refer to Appendix for more details.
Datasets: We considered three popular time-series forecasting datasets: ETTh2, ETTm2 (Zhou et al. 2021) and Weather (Weather 2021). For classification, we considered three real-world datasets, namely, Epilepsy (Andrzejak et al. 2001), EMG (Goldberger et al. 2000a), and Gesture (Liu et al. 2009). We follow the same evaluation setups as proposed in TF-C (Zhang et al. 2022). To simulate varying scenarios of missing values appearing in real-world time-series datasets, we adopt two synthetic masking schemes that we apply on these benchmark datasets, namely missing completely at random (MCAR) masking and periodic masking. Furthermore, we compared our performance on three real-world datasets: PhysioNet-2012 (Silva et al. 2012) for health monitoring, Falling Creek Reservoir (FCR) dataset for modeling lake water quality, and Lake Mendota from the North Temperate Lakes Long-Term Ecological Research program (NTL-LTER; Magnuson, Carpenter, and Stanley 2024) also for modeling lakes. See Appendix for more details.
5 Results and Discussions
5.1 Imputation-based vs. Imputation-free
Figure 6: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline.
(a) Impact of Missing Data Fractions.
@@ 5 Results and Discussions
(b) Impact of the Nature of Missing Values.
@@
Figure 7: Comparison of MSE and F1-score across imputation methods.
(c) Impact of Imputation Methods.
@@
(d) Comparisons on PhysioNet Data. In Figure 7(b), we compare MissTSM with six imputation baselines— M-RNN (Yoon, Zame, and van der Schaar 2018), GP-VAE (Fortuin et al. 2020), BRITS (Cao et al. 2018), Transformer (Vaswani et al. 2017), and SAITS (Du, Côté, and Liu 2023)—on a popular real-world classification dataset, PhysioNet (Silva et al. 2012) following the same evaluation setup as proposed in (Du, Côté, and Liu 2023). MissTSM achieves an impressive F1-score of 57.84%, representing an approximately 15% improvement over the best-performing model (trained on SAITS imputed data). This substantial performance gain on a real-world dataset with missing values highlights the potential of imputation-free or single-stage approaches compared to imputation-based approaches. Please refer to the Appendix for implementation details.
5.2 Comparing Model Agnostic vs. Specialized Models
(a) Comparing MissTSM with GRU-D on Classification.
To analyze the potential of model-agnostic approaches we apply the same MissTSM-integrated MAE model on synthetically masked (80%) classification datasets and compare against GRU-D. From Table 1, we observe that while GRU-D is a specialized model for IMTS data, the proposed model-agnostic still outperforms it significantly. Please refer to the Appendix for more implementation details.
Dataset GRU-D (F1) MissTSM (F1)
@@
Table 1: Comparing MissTSM approach against GRU-D for classification datasets
(b) Comparing MissTSM with Latent ODE on ETTh2.
As discussed above, specialized IMTS models cannot be easily adapted to a different task. To analyze this further, we adapt the Latent ODE model (with ODE-RNN encoder) for a long-term forecasting problem and compare it against our model agnostic approach. We consider a simple setup with 336 context length and 96 prediction length under MCAR masking with varying fractions. From Table 2, we see that Latent ODE struggles to perform long-sequence modeling, with significantly high MSE values. This further motivates the need for model agnostic approaches for handling IMTS data.
(c) Analyzing Model Agnostic Nature of MissTSM.
To further analyze model agnostic capability of the proposed approach we integrate MissTSM with other MTS models like PatchTST and iTransformer. Table 3 shows competitive performance of MissTSM integrated with iTransformer, revealing potential for plugging MissTSM with advanced MTS models for improved performance on downstream tasks even in the presence of missing values with minimal change to the MTS model architecture.
(d) Impact on Real-world Datasets.
We observed that existing benchmark datasets used for forecasting represent a certain level of seasonality which makes it easier for imputation-based models to show adequate performance. However, in many real-world datasets such as those encountered in ecology, there are complex forms of temporal structure in the data beyond simple seasonality. We compare the performance of MissTSM integrated with two MTS models, iTransformer and PatchTST, on two Lake Datasets: Falling Creeks Reservoir and Mendota. Figure 8 shows that MissTSM consistently shows better performance than imputation-based baselines. This further illustrates the importance of imputation-free and model-agnostic approaches for IMTS modeling such as MissTSM.
Fraction Latent ODE (MSE) MissTSM (MSE)
@@
Table 2: Comparing MissTSM with Latent ODE adapted for forecasting.
Dataset
Horizon
Window
iTransformer
+ MissTSM
iTransformer
+ SAITS
iTransformer
+ Spline
ETTh2 96 0.3500.006 0.6450.019 0.3250.013
192 0.4820.027 0.6120.016 0.3400.017
336 0.4040.009 0.5260.015 0.4310.005
720 0.4900.012 0.5050.005 0.4360.017
ETTm2 96 0.2050.001 0.3610.066 0.1760.001
192 0.2680.016 0.4100.064 0.2400.001
336 0.3250.018 0.4580.065 0.2900.002
720 1.0800.188 0.5210.063 0.3850.003
Weather 96 0.1930.002 0.1940.004 0.2020.001
192 0.2400.004 0.2300.002 0.2400.002
336 0.2980.004 0.2800.001 0.2910.001
720 0.3640.003 0.3480.005 0.3570.001
@@
Figure 8: Forecasting performance comparison on Lake datasets across different prediction horizon windows.
6 Conclusions and Future Work
@@
We investigate the performance of existing IMTS models as well as our proposed MissTSM framework on a variety of datasets and tasks with varying conditions of missing values. We show that imputation-based frameworks built on simple imputations such as Spline and SAITS perform well when the amount of missingness is small or there is periodic structure in the data (e.g., in Weather data) that is easy to approximate. However, imputation-based approaches show poor performance at larger missing value fractions and when missing values have limited periodic patterns (e.g., on the lake datasets). We also show that MissTSM, which is an imputation-free and model-agnostic framework shows competitive performance across most datasets, tasks, and settings compared to imputation-based and existing imputation-free specialized models. Our proposed MissTSM also has limitations. For example, a limitation of the MFAA layer is that it does not learn the non-linear temporal dynamics and relies on the subsequent transformer encoder blocks to learn the dynamics. Future work can explore modifications of the MFAA layer such that it can jointly learn the cross-channel correlations with the non-linear temporal dynamics.
@@ A.1 Forecasting
ETT. The ETT (Zhou et al. 2021) dataset captures load and oil temperature data from electricity transformers. ETTh2 includes 17,420 hourly observations, while ETTm2 comprises 69,680 15-minute observations. Both datasets span two years and contain 7 variates each.
Weather. Weather (Weather 2021) is a 10-minute frequency time-series dataset recorded throughout the year 2020 and consists of 21 meteorological indicators, like humidity, air temperature, etc.
Solar-Energy. Solar-energy (Lai et al. 2018) dataset captures solar power generation data from 137 photovoltaic (PV) plants in the year 2006, with data sampled at a resolution of every 10 min, providing rich MTS data. We report results on Solar-Energy dataset in the appendix only, as supplementary forecasting evaluation.
Following (Wu et al. 2021) and (Lai et al. 2018), we use a train-validation-test split of 6:2:2 for the ETT and Solar-Energy datasets and 7:1:2 for the Weather dataset.
Ecology Dataset - Falling Creek Reservoir (FCR). The FCR dataset is generated by combining data from the following data packages (Carey et al. 2023, 2024; Carey and Breef-Pilz 2023b, a; Carey, Breef-Pilz, and Woelmer 2023), published in the Environmental Data Initiative repository. The dataset contains daily median meteorological and water quality observations from FCR (Virginia, USA), collected over a period of 3 years spanning from 2018-08-01 to 2021-12-22. As part of the pre-processing, we removed variables with more than 90% missing values. Specifically, two columns were excluded with the remaining dataset consisting of 1240 time points with an overall missing rate of 53.12%.
Ecology Dataset - Mendota. The Mendota dataset contains daily averaged meteorological and water quality observations from Lake Mendota (Dane County, WI, USA). Data were collected using an instrumented buoy at the surface of the water during the ice-free season from July 2006 through November 2023 (Magnuson, Carpenter, and Stanley 2024). The data were cleaned by filtering out flagged values before use. We applied preprocessing and removed variables with more than 90% missing data. This resulted in dropping three columns with 99% missing values, and one column with 98.1% missing values. The resulting dataset contains 6321 time points, with an overall missing rate of 50.07%.

### v3

@@ 5 Results and Discussions
Figure 4: Performance comparison against different TS Baselines imputed with SAITS, across different missing data fractions.
@@
Figure 5: Classification F1 scores on three datasets, EMG, Epilepsy, and Gesture. Masking fractions considered: 0.2, 0.4, 0.6, 0.8.
@@
Figure 6: Comparison of different masking methods (70% missing fraction): MCAR vs. Periodic Masking for ETTh2, ETTm2, and Weather datasets. SA stands for SAITS and SP stands for Spline. Lower MSE is better
5.1.2 Impact of the Nature of Missing Values.
@@
Figure 7: Comparison of MSE (lower is better) and F1-score across imputation methods.
@@
In Figure 7(b), we compare MissTSM with six imputation baselines - M-RNN (Yoon et al., 2018b), GP-VAE (Fortuin et al., 2020), BRITS (Cao et al., 2018), Transformer (Vaswani et al., 2017), and SAITS (Du et al., 2023) - on a popular real-world classification dataset, PhysioNet (Silva et al., 2012) following the same evaluation setup as proposed in (Du et al., 2023). MissTSM achieves an impressive F1-score of 57.84%, representing an approximately 15% improvement over the best-performing model (trained on SAITS imputed data). This substantial performance gain on a real-world dataset with missing values highlights the potential of imputation-free or single-stage approaches compared to imputation-based approaches.
5.2 Comparing Model-Agnostic vs. Specialized Models
5.2.1 Analyzing MissTSM on IMTS Classification and Forecasting
We evaluate MissTSM on both classification and forecasting tasks for irregular multivariate time series. To illustrate the generality of our approach, we study two case models: (i) GRU-D, a specialized classifier for irregularly sampled data, and (ii) Latent ODE, a continuous-time generative model not originally designed for forecasting but adapted here to a long-term prediction setting. These case studies emphasize how specialized methods struggle when moved beyond their intended use, underscoring the value of model-agnostic approaches.
IMTS Classification. We conduct experiments on the IMTS classification task using the P12 (Goldberger et al., 2000b) and P19 (Reyna et al., 2020) datasets, following the same evaluation protocol as Luo et al. (2025). We report the baseline results for the considered models directly from Luo et al. (2025). Table 1 highlights the strong potential of model-agnostic approaches; integrating the MissTSM layer, can achieve performance on par with or exceeding that of several well-known IMTS models.
Methods P19 P12
AUROC AUPRC AUROC AUPRC
GRU-D 88.7 1.2 56.2 2.3 79.6 0.6 41.7 1.8
ODE-RNN 87.1 1.0 52.6 3.2 78.8 0.6 37.4 2.6
SeFT 84.0 0.3 49.3 0.5 78.1 0.5 35.9 0.8
mTAND 82.9 0.9 32.2 1.5 85.3 0.3 49.3 1.0
Raindrop 87.6 2.7 61.1 1.4 82.0 0.6 42.7 1.7
MTGNN 88.5 1.0 55.8 1.5 82.1 1.5 41.8 2.1
MissTSM 88.8 1.3 56.5 1.2 82.2 0.5 43.8 1.1
Table 2: Comparing (F1 scores) MissTSM approach against GRU-D for classification datasets.
Dataset GRU-D MissTSM
@@ 5.2.1 Analyzing MissTSM on IMTS Classification and Forecasting
Table 3: Comparing (MSE values) MissTSM with Latent ODE adapted for forecasting
Fraction Latent ODE MissTSM
@@
Comparing MissTSM with GRU-D on Classification. To analyze the potential of model-agnostic approaches we apply the same MissTSM-integrated MAE model on synthetically masked (80%) classification datasets and compare against GRU-D. From Table 3, we observe that while GRU-D is a specialized model for IMTS data, the proposed model-agnostic still outperforms it significantly. Please refer to the Appendix for more implementation details.
Comparing MissTSM with Latent ODE on ETTh2. As discussed above, specialized IMTS models cannot be easily adapted to a different task. To analyze this further, we adapt the Latent ODE model (with ODE-RNN encoder) for a long-term forecasting problem and compare it against our model-agnostic approach. We consider a simple setup with 336 context length and 96 prediction length under MCAR masking with varying fractions. From Table 3, we see that Latent ODE struggles to perform long-sequence modeling, with significantly high MSE values. Moreover, ODE-based methods incur considerable computational costs, which grow even more pronounced for long-term modeling.
5.2.2 Analyzing Model-Agnostic Nature of MissTSM.
To further analyze model-agnostic capability of the proposed approach we integrate MissTSM with other MTS models like PatchTST and iTransformer. Tables 5 and 5 show competitive performance of MissTSM integrated with PatchTST, revealing potential for plugging MissTSM with advanced MTS models for improved performance on downstream tasks even in the presence of missing values with minimal change to the MTS model architecture. Please refer to appendix for additional results.
Table 4: MSE (meanstd) for PatchTST with MissTSM under 60% masking.
Dataset Horizon Window PatchTST + MissTSM PatchTST + SAITS PatchTST + Spline
ETTh2 96 0.3170.004 0.5030.013 0.3240.013
192 0.3770.009 0.5120.011 0.3990.017
336 0.3800.011 0.4100.012 0.4310.005
720 0.5140.033 0.4110.002 0.4360.017
ETTm2 96 0.2020.005 0.3220.045 0.1690.000
192 0.2610.002 0.3590.036 0.2270.000
336 0.3130.001 0.4080.043 0.2850.001
720 0.4200.027 0.4590.035 0.3760.001
Weather 96 0.2060.014 0.1690.001 0.2700.110
192 0.2760.027 0.2120.000 0.2870.080
336 0.3090.024 0.2630.001 0.3250.065
720 0.3400.003 0.3330.001 0.3910.060
Table 5: MSE (meanstd) for PatchTST with MissTSM under 70% masking.
Dataset Horizon Window PatchTST + MissTSM PatchTST + SAITS PatchTST + Spline
ETTh2 96 0.3220.004 0.5480.050 0.3170.009
192 0.3820.011 0.5610.056 0.3800.005
336 0.3840.008 0.4680.059 0.3720.007
720 0.6210.025 0.4970.085 0.4190.015
ETTm2 96 0.2130.006 0.4050.079 0.1770.009
192 0.2660.003 0.4470.086 0.2360.009
336 0.3150.004 0.4940.095 0.2930.007
720 0.4320.025 0.5290.092 0.3860.009
Weather 96 0.2040.014 0.1660.013 0.2620.122
192 0.2490.030 0.2070.009 0.2790.092
336 0.3040.026 0.2570.007 0.3170.075
720 0.3580.012 0.3290.008 0.3830.071
5.2.3 Impact on Real-world Datasets
@@
Figure 8: Forecasting performance comparison on Lake datasets across different prediction horizon windows. Lower MSE (y-axis) is better
We observed that existing benchmark datasets used for forecasting represent a certain level of seasonality which makes it easier for imputation-based models to show adequate performance. However, in many real-world datasets such as those encountered in ecology, there are complex forms of temporal structure in the data beyond simple seasonality. We compare the performance of MissTSM integrated with two MTS models, iTransformer and PatchTST, on two Lake Datasets: Falling Creeks Reservoir and Mendota. Figure 8 reports masked MSE - MSE computed only on observed points - comparing MissTSM against imputation-based baselines. Competitive performance shown by MissTSM on both the real-world missing datasets further motivates the idea of imputation-free and model-agnostic approaches for IMTS modeling.
5.3 Ablations
In the ablation experiments, we evaluate the impact of integrating the MissTSM layer. We compare MAE with MissTSM against standard MAE (without MissTSM), using spline and SAITS imputation as additional baselines. The goal here is to understand the additional value of adding the MissTSM layer instead of modeling on imputed data. For forecasting (Fig. 9) and classification (Fig. 10), MissTSM consistently improves performance. In forecasting, MissTSM-MAE outperforms all MAE variants, while in classification, it is consistently comparable or superior across all three datasets.
6 Conclusion
@@
We investigate the performance of existing IMTS models as well as our proposed MissTSM framework on a variety of datasets and tasks with varying conditions of missing values. We show that imputation-based frameworks built on simple imputations perform well when the amount of missingness is small or there is periodic structure in the data (e.g., in Weather data) that is easy to approximate. However, imputation-based approaches show poor performance at larger missing value fractions and when missing values have limited periodic patterns (e.g., on the lake datasets). We also show that MissTSM, which is an imputation-free and model-agnostic framework, demonstrates competitive performance across most datasets, tasks, and settings compared to imputation-based and existing imputation-free specialized models. We hope our findings could inspire further research into developing flexible, model-agnostic adapters for handling the challenges in irregularly-sampled time-series data.
Limitations and Future Directions. (1) A limitation of the MFAA layer is that it doesn’t learn the non-linear temporal dynamics and relies on the subsequent transformer encoder blocks to learn the dynamics. Future work can explore modifications of the MFAA layer such that it can jointly learn the cross-channel correlations with the non-linear temporal dynamics. (2) Independent embedding of each time-feature token can become computationally expensive in high-dimensional multivariate systems.
Broader Impact Statement
Validation requirements and safety considerations when used in healthcare sectors MissTSM is a decision-support tool and would require rigorous clinical validation to ensure reliable predictions. We would recommend that the model must be used in an "expert-in-the-loop" system
Guidance on when the method is reliable vs when it may fail The method is reliable in challenging scenarios with low periodicity, high missingness and high number of variates. On low variate, highly periodic data, a simpler approach like Spline + Backbone can be more effective
The potential issue of over-trust in imputation-free outputs We agree that this is a critical concern. Our recommendation would be to pair MissTSM’s predictions with a reliable uncertainty quantification metric. This metric should reflect the sparsity (or quality) of the input data, warning the user when the predictions are of low confidence.
Acknowledgments
This work was supported in part by NSF awards IIS-2239328 and DEB-2213550. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-public-access-plan).
@@ A.1 Forecasting
ETT. The ETT (Zhou et al., 2021) dataset captures load and oil temperature data from electricity transformers. ETTh2 includes 17,420 hourly observations, while ETTm2 comprises 69,680 15-minute observations. Both datasets span two years and contain 7 variates each.
Weather. Weather (Weather, 2021) is a 10-minute frequency time-series dataset recorded throughout the year 2020 and consists of 21 meteorological indicators, like humidity, air temperature, etc.
Solar-Energy. Solar-energy (Lai et al., 2018) dataset captures solar power generation data from 137 photovoltaic (PV) plants in the year 2006, with data sampled at a resolution of every 10 min, providing rich MTS data. We report results on Solar-Energy dataset in the appendix only, as supplementary forecasting evaluation.
Following Wu et al. (2021) and Lai et al. (2018), we use a train-validation-test split of 6:2:2 for the ETT and Solar-Energy datasets and 7:1:2 for the Weather dataset.
Ecology Dataset - Falling Creek Reservoir (FCR). The FCR dataset is generated by combining data from the following data packages, Carey et al. (2023a; 2024); Carey and Breef-Pilz (2023b; a); Carey et al. (2023b), published in the Environmental Data Initiative repository. The dataset contains daily median meteorological and water quality observations from FCR (Virginia, USA), collected over a period of 3 years spanning from 2018-08-01 to 2021-12-22. As part of the pre-processing, we removed variables with more than 90% missing values. Specifically, two columns were excluded with the remaining dataset consisting of 1240 time points with an overall missing rate of 53.12%.
Ecology Dataset - Mendota. The Mendota dataset contains daily averaged meteorological and water quality observations from Lake Mendota (Dane County, WI, USA). Data were collected using an instrumented buoy at the surface of the water during the ice-free season from July 2006 through November 2023 (Magnuson et al., 2024). The data were cleaned by filtering out flagged values before use. We applied preprocessing and removed variables with more than 90% missing data. This resulted in dropping three columns with 99% missing values, and one column with 98.1% missing values. The resulting dataset contains 6321 time points, with an overall missing rate of 50.07%.